OpenStudy (anonymous):

Find the perimeter of ABC. Round your answer to the nearest tenth. A (-4,-1) B (4,5) C (4,-2) The distance formula is (x2-x1)^2 + (y2-y1)^2

jimthompson5910 (jim_thompson5910):

it's actually d = sqrt( (x2-x1)^2 + (y2-y1)^2 ) but you are close

jimthompson5910 (jim_thompson5910):

you use that formula to find the lengths of AB AC BC then add up these lengths

OpenStudy (anonymous):

i'm still confused about how to do it, but thanks

jimthompson5910 (jim_thompson5910):

I'll show you how to find the distance from A to B to get the length AB

OpenStudy (e.mccormick):

A perimiter is the distance of all three sides. The points you have ben given are where the sides meet.

jimthompson5910 (jim_thompson5910):

A = (-4,-1) B = (4,5) (x1,y1) = (-4, -1) ----> x1 = -4, y1 = -1 (x2,y2) = (4, 5) ----> x2 = 4, y2 = 5 ------------------------------------------------------- d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((4-(-4))^2+(5-(-1))^2) d = sqrt((4+4)^2+(5+1)^2) d = sqrt((8)^2+(6)^2) d = sqrt(64+36) d = sqrt(100) d = 10 So exact distance between the two points, A and B, is 10 units. So AB is 10 units long

jimthompson5910 (jim_thompson5910):

do the same for AC and BC then add up the 3 sides

jimthompson5910 (jim_thompson5910):

Does that help?

OpenStudy (e.mccormick):